# NON-CLASSICAL GENERATING SETS FOR FUCHSIAN SCHOTTKY GROUPS

### v6

@@ 2. THE KEY LEMMAS
$x=-2,-\frac{5}{3},-1,$ and $-\frac{2}{3}$.
Now, let $\lambda=\lvert x\rvert,$ where $\lvert x\rvert$ denotes the modulus value of the numbers $-2$ and $-\frac{5}{3}$, i.e., $\lambda=2$ and $1.6666666667$ (approx.) $(>1)$.
We have supposed $x$ to be a positive real number greater than $1$ but the equation $09x^{4}+48x^{3}+91x^{2}+72x+20=$ 0 could not give such a real number as a root. On the other hand, if we want to use the technique from Yamamoto’s paper [11], we have to take $209x^{4}+48x^{3}+91x^{2}+72x+$ 20 equal to zero. Instead, we could not find our required sufficiently small positive delta. Consequently, for simplicity, we have taken the modulus values $(>1)$ of the roots of that equation without loss of generality.
@@
[…]
@@
Now, we aim to show that the length of any image of ${\Psi}$ under $\beta_{2t}$ is less than the length of ${\Psi}$, which we have established by proving that $\lvert\beta^{\prime}_{2t}(x)\rvert<1,$ $\forall$ $x$in ${\Psi}$. In fact, we have succeeded by utilizing the method of induction concerning $t$.
@@
[…]
@@
$(i)$ Case I, for $s_{1}=0:$ In the first case, our goal is to prove that |$\lvert\{({h^{*}}_{(3,2)_{\kappa}})^{s_{2}}\}^{\prime}(x)\rvert<1$ $\forall$ $x$in ${\Psi}$ by using the method of induction on $s_{2}$. Now if we observe the case $s_{2}>0$, for $s_{2}=1$ the induction becomes trivial. For $s_{2}=2$, we obtain the following.
@@
[…]
@@
the position of $\psi_{1}$ will be on the right of this domain whereas $\psi_{2}$ situated to the left of that region, and $\psi_{3}$ is located to the right of this zone. Further, for some very small $\kappa$, $\psi_{4}$ is also outside that range. Hence, we obtain $\forall$ $x$in $\Psi$, |$\lvert\{({h^{*}}_{(3,2)_{\kappa}})^{s_{2}}\}^{\prime}(x)\rvert<1$. Now, our goal is to prove that the condition |$\lvert\{({h^{*}}_{(3,2)_{\kappa}})^{s_{2}}\}^{\prime}(x)\rvert<1$ is true for $s_{2}=n+1$ when it happens for $s_{2}=n$.
@@
The situation is very similar for the case when $s_{2}<0$. First, we show it for $s_{2}=-2$, after that, our target will be to prove that assuming |$\lvert\{({h^{*}}_{(3,2)_{\kappa}})^{s_{2}}\}^{\prime}(x)\rvert<1$ is true for $s_{2}=n$ (where, $n<0)$, then it is also true for $s_{2}=n-1$. In this manner for $s_{2}<0$ one can follow the exact method as we have discussed in the previous case. Therefore, Case I is done.
@@
$\lvert(\beta_{2})^{\prime}(x)\rvert$ $=$ |$\lvert\{({h^{*}}_{(3,2)_{\kappa}})^{s_{2}}({h^{**}}_{(4,1)_{\kappa}})^{s_{1}}\}^{\prime}(x)\rvert$
@@
by using Case I when $s_{1}=0$ and applying all the regions (for $(\text{for }s_{1}<0)$ that are outside the semi-circle $SC_{3,\kappa}$.
@@
$(iii)$ Case III, for $s_{1}>0:$ For the first step of the induction concerning $t$, we aim to derive that for $s_{1}>0$, the image of ${\Psi}$ under the transformations $\{({h^{*}}_{(3,2)_{\kappa}})^{s_{2}}({h^{**}}_{(4,1)_{\kappa}})^{s_{1}}\}$ is less than ${\Psi}$. Now, for the subcases, we first examine the case for $s_{2}>0$. Observe that, if $s_{1}=1$ then from our assumption, we get that the transformation $({h^{**}}_{(4,1)_{\kappa}})$ sends ${\psi}_{3}$ to ${\psi}_{2}$ and ${\psi}_{4}$ to ${\psi}_{1}$, and also we already have |$\lvert\{({h^{*}}_{(3,2)_{\kappa}})^{s_{2}}\}^{\prime}(x)\rvert<1$ $\forall$ $x$in ${\Psi}$. Further, when $s_{1}>1$, the images of ${\psi}_{1}$ and ${\psi}_{2}$ under the transformation $({h^{**}}_{(4,1)_{\kappa}})^{s_{1}}$ are outside the semi-circle $SC_{3,\kappa}$. Again, outside the semi-circle $SC_{3,\kappa}$, the Möbius transformation $({h^{*}}_{(3,2)_{\kappa}})$ is contracting, since we get
@@
[…]
@@
$<$ $\frac{1}{\lvert x+(\lambda+2)\rvert^{s_{2t+1}}}\times\Big{\lvert}\frac{d\{({h^{*}}_{(3,2)_{\kappa}})^{s_{2t+2}}(y)\}}{dy}\Big{\rvert},\text{ replacing }y\text{ by }\{({h^{**}}_{(4,1)_{\kappa}})^{s_{2t+1}}\beta_{2t}(x)\}.$
[…]
@@
Subcase II(b): In this case, the given condition is $(\lambda-1)\leq\lvert\beta_{2n}(x)\rvert<{e}$. So, this condition yields that $\lvert({h^{**}}_{(4,1)_{\kappa}})^{s_{2t-1}}\beta_{2t-2}(x)\rvert$ $<(\lambda-1)$. Therefore, $s_{2t-1}<0$. Now, one can easily compute the value of $\lvert\beta^{\prime}_{2t-2}(x)\rvert$ by using the technique operated to evaluate the value of $\lvert\beta^{\prime}_{2t+2}(x)\rvert$ in the subcase II(a). Then, the proof of the subcase II(b) follows. So, we have established that $\lvert\beta^{\prime}_{2t+2}(x)\rvert<1$. Consequently, we have finished the induction.
@@
So, when $z$ $<$ $\frac{\lambda^{2}+5\lambda+5}{\lambda+3}$ (for $\kappa\rightarrow 0^{+}$), the value of the transformation $\Big{(}{h^{**}}_{(4,1)_{\kappa}}\Big{)}^{-1}$ is less than $1$.
@@
Note that, $\Big{(}{h^{**}}_{(4,1)_{\kappa}}\Big{)}^{2}(z)$ $=$ $h(z)$,
[…]
where, $h:z\rightarrow(\lambda+2)z$, where $z$in $\mathbb{H}^{2}$.
In fact, for simplicity, we want to go forward to the function $h$ instead of the transformation $\Big{(}{h^{**}}_{(4,1)_{\kappa}}\Big{)}^{2}$, without loss of generality. Now, if we utilize the transformation ${h^{**}}_{(4,1)_{\kappa}}$ twice to the domain mentioned in Lemma 2.2, i.e., the fragment $[-(\lambda+3),-(\lambda+1)]\cup[(\lambda+1),(\lambda+3)]$, the component $\{\Omega\textbf{(}{h^{**}}_{(4,1)_{\kappa}}\Gamma_{S_{\kappa}}^{N}({h^{**}}_{(4,1)_{\kappa}})^{-1}\textbf{)}\cap i\mathbb{R}^{+}\}$ will touch the vertical segment $i[1,$ $(\lambda+2)(\lambda+3)]$. Since the transformation $\Big{(}{h^{**}}_{(4,1)_{\kappa}}\Big{)}^{-1}$ is contracting, we can use the function $\Big{(}{h^{**}}_{(4,1)_{\kappa}}\Big{)}^{-1}$ several times to extend the standing piece $i[1,$ $(\lambda+2)(\lambda+3)]$ to the whole upright portion, i.e., $i[0,$ $(\lambda+2)(\lambda+3)]$. Observe that, this act doesn’t affect the size of the interval. So the lemma follows. ∎
@@
$\lvert Y^{\prime}_{k_{i}}-P\rvert+\lvert Y^{\prime}_{k_{j}}-P\rvert$$\geq$ $\lvert Y^{\prime}_{k_{i}}-Y^{\prime}_{k_{j}}\rvert$ (5)
@@
It is known that the fixed points of the transformation ${h}_{(4,1)_{\kappa}}^{**}$ are $\tau$ and $-\tau$. Now, since the semi-circle ${SC}_{\kappa}$ separates the fixed points of the transformation ${h}_{(4,1)_{\kappa}}^{**}$, any image of the semi-circular curve ${SC}_{\kappa}$ under the map ${h}_{(4,1)_{\kappa}}^{**}$ will also separate these fixed points. These fixed points are situated in either one of the curves in the set ${SC}^{*}_{\kappa}=\{{SC}^{1}_{\kappa},{SC}^{2}_{\kappa},{SC}^{3}_{\kappa},...,{SC}^{\textbf{I}^{+}}_{\kappa}\}$ or outside the semi-circle $SC_{1,\kappa}$. The image can not be inside ${SC}_{\kappa}^{j}$ (where $j=1,2,3,...,\textbf{I}^{+}$). Also, the image can not be the semi-circle ${SC}_{\kappa}^{j}$ either, since ${SC}_{\kappa}^{j}$ is tangential to ${SC}_{\kappa}$ at the point $P$. Therefore it must be outside of the semi-circle ${SC}_{\kappa}^{j}$. So, the images of the $Y_{k}^{\prime}$ for the map ${h}_{(4,1)_{\kappa}}^{**}$ must be outside the $Y_{k}^{\prime}$ themselves.
@@
[…]
@@
Since the semi-circular curves ${SC}_{\kappa}^{j}$ and ${SC}_{\kappa}^{j+1}$, along with two other curves in the set $\{q(SC_{i,{\kappa}}):q\in\Gamma^{{N}}_{S_{\kappa}},$ $i=1,2,3,\text{ and }4\}$ bound a fundamental domain for $\Gamma_{S_{\kappa}}^{{N}}$, where $\Gamma_{S_{\kappa}}^{{N}}$ $=$ $<{h^{*}}_{(3,2)_{\kappa}},{h^{**}}_{(4,1)_{\kappa}}>$, we assert that ${h^{**}}_{(4,1)_{\kappa}}(SC_{\kappa}^{j})\cap SC_{\kappa}^{j+1}=\phi$. Let, $M$ and $N$ be two points lying on the semi-circle ${SC}_{\kappa}^{j}$ (where $j=1,2,3,...,\textbf{I}^{+}$). Note that the transformation ${h^{**}}_{(4,1)_{\kappa}}$ preserves any line through the points ±$\pm\tau$. Now, if we take a line ${H}_{(4,1)_{\kappa}}$ through the point $\{-\tau\}$ and the point $({x}_{j},0)$ we have the distance between the point $M$ and ${C}_{j}$ is $\{{r}_{j}-\textbf{(}\lvert{C}_{j}\rvert+\tau\textbf{)}\}$, which is away from the fixed point $\{-\tau\}$. Further, the distance between the point $N$ and ${C}_{j}$ is $\{{r}_{j}+\textbf{(}\lvert{C}_{j}\rvert+\tau\textbf{)}\}$. Also, the image of point $M$ under the transformation ${h^{**}}_{(4,1)_{\kappa}}$ situates on the line ${H}_{(4,1)_{\kappa}}$ and further from the point $\{-\tau\}$ than the point $N$.
@@
$\lvert{x}_{j+1}+\tau\rvert$$\leq$ $\{\lvert{x}_{j+1}\rvert+\lvert\tau\rvert\}$ (10)

### v7

@@ 3. PROOF OF THE LEMMAS
$x=-2,-\frac{5}{3},-1,$ and $\frac{2}{3}$.
Now, let $\lambda=\lvert x\rvert$, where $\lvert x\rvert$ denotes the modulus value of the numbers 2-2 and $\frac{5}{3}$, i.e., $\lambda=2$ and $1.6666666667$ (approx.) $(>1)$.
We have supposed $x$ to be a positive real number greater than $1$ but the equation $9x^{4}+48x^{3}+91x^{2}+72x+20=0$ could not give such a real number as a root. On the other hand, if we want to use the technique from Yamamoto’s paper [11], we have to take $9x^{4}+48x^{3}+91x^{2}+72x+20$ equal to zero. Instead, we could not find our required sufficiently small positive delta. Consequently, for simplicity, we have taken the modulus values $(>1)$ of the roots of that equation without loss of generality.
@@
$\text{ $l({\Psi}$})$ $<$ $\Big\{\frac{1}{(4x^{2}+9x+4-2x\kappa+x\kappa^{2})}\times\sqrt{2.01}\times\sqrt{\kappa}\times\sqrt{6\lambda^{4}+16\lambda^{3}+3\lambda^{2}-16\lambda-8}\Big\}$
@@
Now, we aim to show that the length of any image of $\Psi$} under $\beta_{2t}$ is less than the length of $\Psi$}, which we have established by proving that $\lvert\beta^{\prime}_{2t}(x)\rvert<1,$ $\forall$ $x$ $\in$ $\Psi$}. In fact, we have succeeded by utilizing the method of induction concerning $t$.
@@
$\text{ and }\lvert({h^{**}}_{(4,1)_{\kappa}})^{\prime}(x)\rvert$ $=$ $\Big\lvert\frac{(1-\kappa)^{2}}{\{x+(\lambda+2)\}^{2}}\Big\rvert$.
@@
$(i)$ Case I, for $s_{1}=0:$ In the first case, our goal is to prove that $\lvert\{({h^{*}}_{(3,2)_{\kappa}})^{s_{2}}\}^{\prime}(x)\rvert<1$ $\forall$ $x$ $\in$ $\Psi$} by using the method of induction on $s_{2}$. Now if we observe the case $s_{2}>0$, for $s_{2}=1$ the induction becomes trivial. For $s_{2}=2$, we obtain the following.
@@
$\text{ and }x$ $>$ $\frac{1}{4}\{(-2\lambda+1)+\sqrt{(2\lambda+1)^{2}+16(1-\kappa)^{2}}\},$
@@
the position of $\psi_{1}$ will be on the right of this domain whereas $\psi_{2}$ situated to the left of that region, and $\psi_{3}$ is located to the right of this zone. Further, for some very small $\kappa$, $\psi_{4}$ is also outside that range. Hence, we obtain $\forall$ $x$ $\in$ $\Psi$, $\lvert\{({h^{*}}_{(3,2)_{\kappa}})^{s_{2}}\}^{\prime}(x)\rvert<1$. Now, our goal is to prove that the condition $\lvert\{({h^{*}}_{(3,2)_{\kappa}})^{s_{2}}\}^{\prime}(x)\rvert<1$ is true for $s_{2}=n+1$ when it happens for $s_{2}=n$.
@@
The situation is very similar for the case when $s_{2}<0$. First, we show it for $s_{2}=-2$, after that, our target will be to prove that assuming $\lvert\{({h^{*}}_{(3,2)_{\kappa}})^{s_{2}}\}^{\prime}(x)\rvert<1$ is true for $s_{2}=n$ (where, $n<0)$, then it is also true for $s_{2}=n-1$. In this manner for $s_{2}<0$ one can follow the exact method as we have discussed in the previous case. Therefore, Case I is done.
@@
$\lvert(\beta_{2})^{\prime}(x)\rvert$ $=$ $\lvert\{({h^{*}}_{(3,2)_{\kappa}})^{s_{2}}({h^{**}}_{(4,1)_{\kappa}})^{s_{1}}\}^{\prime}(x)\rvert$
@@
[…]
@@
$(iii)$ Case III, for $s_{1}>0:$ For the first step of the induction concerning $t$, we aim to derive that for $s_{1}>0$, the image of $\Psi$} under the transformations $\{({h^{*}}_{(3,2)_{\kappa}})^{s_{2}}({h^{**}}_{(4,1)_{\kappa}})^{s_{1}}\}$ is less than $\Psi$}. Now, for the subcases, we first examine the case for $s_{2}>0$. Observe that, if $s_{1}=1$ then from our assumption, we get that the transformation $({h^{**}}_{(4,1)_{\kappa}})$ sends ${\psi}_{3}$ to ${\psi}_{2}$ and ${\psi}_{4}$ to ${\psi}_{1}$, and also we already have $\lvert\{({h^{*}}_{(3,2)_{\kappa}})^{s_{2}}\}^{\prime}(x)\rvert<1$ $\forall$ $x$ $\in$ $\Psi$}. Further, when $s_{1}>1$, the images of ${\psi}_{1}$ and ${\psi}_{2}$ under the transformation $({h^{**}}_{(4,1)_{\kappa}})^{s_{1}}$ are outside the semi-circle $SC_{3,\kappa}$. Again, outside the semi-circle $SC_{3,\kappa}$, the Möbius transformation $({h^{*}}_{(3,2)_{\kappa}})$ is contracting, since we get
@@
$\text{ and }B_{2}$ $=$ $\Big\lvert({h^{*}}_{(3,2)_{\kappa}})\Big\{\Big(\frac{(\lambda+2){\psi}_{2a}+A}{{\psi}_{2a}+(\lambda+2)}\Big)^{s_{1}}\Big\}-({h^{*}}_{(3,2)_{\kappa}})\Big\{\Big(\frac{(\lambda+2){\psi}_{2b}+A}{{\psi}_{2b}+(\lambda+2)}\Big)^{s_{1}}\Big\}\Big\rvert$
@@
$<$ $\frac{1}{\lvert x+(\lambda+2)\rvert^{s_{2t+1}}}\times\Big\lvert\frac{d\{({h^{*}}_{(3,2)_{\kappa}})^{s_{2t+2}}(y)\}}{dy}\Big\rvert,\text{ replacing }y\text{ by }\{({h^{**}}_{(4,1)_{\kappa}})^{s_{2t+1}}\beta_{2t}(x)\}$
$<$ $\frac{1}{\lvert x+(\lambda+2)\rvert^{s_{2t+1}}}\times\Big\lvert\frac{(1-\kappa)^{2}}{\Big\{\Big(\frac{(1-\kappa)^{2}}{\{x+(\lambda+2)\}}\Big)^{s_{2t+1}}+\lambda\Big\}^{2}}\Big\rvert$
@@
Subcase II(b): In this case, the given condition is $(\lambda-1)\leq\lvert\beta_{2n}(x)\rvert<{e}$. So, this condition yields that $\lvert({h^{**}}_{(4,1)_{\kappa}})^{s_{2t-1}}\beta_{2t-2}(x)\rvert$ $(\lambda-1)$. Therefore, $s_{2t-1}<0$. Now, one can easily compute the value of $\lvert\beta^{\prime}_{2t-2}(x)\rvert$ by using the technique operated to evaluate the value of $\lvert\beta^{\prime}_{2t+2}(x)\rvert$ in the subcase II(a). Then, the proof of the subcase II(b) follows. So, we have established that $\lvert\beta^{\prime}_{2t+2}(x)\rvert<1$. Consequently, we have finished the induction.
@@
So, when $z$ << $\frac{\lambda^{2}+5\lambda+5}{\lambda+3}$ (for $\kappa\rightarrow 0^{+}$), the value of the transformation $\Big({h^{**}}_{(4,1)_{\kappa}}\Big)^{-1}$ is less than $1$.
@@
Note that, $\Big({h^{**}}_{(4,1)_{\kappa}}\Big)^{2}(z)$ == $h(z)$,
$\text{ for }z=\frac{-(2\lambda^{3}+10\lambda^{2}-15\lambda+7)\pm\sqrt{4\lambda^{6}+56\lambda^{5}+168\lambda^{4}+104\lambda^{3}+885\lambda^{2}+1042\lambda+433}}{4(\lambda+2)^{2}},$
where, $h:z\rightarrow(\lambda+2)z$, where $z$ $\in$ $\mathbb{H}^{2}$.
In fact, for simplicity, we want to go forward to the function $h$ instead of the transformation $\Big({h^{**}}_{(4,1)_{\kappa}}\Big)^{2}$, without loss of generality. Now, if we utilize the transformation ${h^{**}}_{(4,1)_{\kappa}}$ twice to the domain mentioned in Lemma 3.2, i.e., the fragment $(\lambda+3),-(\lambda+1)]\cup[(\lambda+1),(\lambda+3)]$, the component $\{\Omega\textbf{(}{h^{**}}_{(4,1)_{\kappa}}\Gamma_{S_{\kappa}}^{N}({h^{**}}_{(4,1)_{\kappa}})^{-1}\textbf{)}\cap i\mathbb{R}^{+}\}$ will touch the vertical segment $i[1,$ $(\lambda+2)(\lambda+3)]$. Since the transformation $\Big({h^{**}}_{(4,1)_{\kappa}}\Big)^{-1}$ is contracting, we can use the function $\Big({h^{**}}_{(4,1)_{\kappa}}\Big)^{-1}$ several times to extend the standing piece $i[1,$ $(\lambda+2)(\lambda+3)]$ to the whole upright portion, i.e., $i[0,$ $(\lambda+2)(\lambda+3)]$. Observe that, this act doesn’t affect the size of the interval. So the lemma follows. ∎
@@
[…]
@@
It is known that the fixed points of the transformation ${h}_{(4,1)_{\kappa}}^{**}$ are $\tau$ and $\tau$. Now, since the semi-circle ${SC}_{\kappa}$ separates the fixed points of the transformation ${h}_{(4,1)_{\kappa}}^{**}$, any image of the semi-circular curve ${SC}_{\kappa}$ under the map ${h}_{(4,1)_{\kappa}}^{**}$ will also separate these fixed points. These fixed points are situated in either one of the curves in the set ${SC}^{*}_{\kappa}=\{{SC}^{1}_{\kappa},{SC}^{2}_{\kappa},{SC}^{3}_{\kappa},...,{SC}^{\textbf{I}^{+}}_{\kappa}\}$ or outside the semi-circle $SC_{1,\kappa}$. The image can not be inside ${SC}_{\kappa}^{j}$ (where $j=1,2,3,...,\textbf{I}^{+}$). Also, the image can not be the semi-circle ${SC}_{\kappa}^{j}$ either, since ${SC}_{\kappa}^{j}$ is tangential to ${SC}_{\kappa}$ at the point $P$. Therefore it must be outside of the semi-circle ${SC}_{\kappa}^{j}$. So, the images of the $Y_{k}^{\prime}$ for the map ${h}_{(4,1)_{\kappa}}^{**}$ must be outside the $Y_{k}^{\prime}$ themselves.
@@
$<$ $\{(\lambda+2)-(-\tau)\}\lvert Y_{k}^{\prime}\rvert\text{ (see, Section 2 for details) }$
$\text{ i.e., }\lvert Y_{k+1}^{\prime}\rvert$ $<$ $(\lambda+2+\tau)\lvert Y_{k}^{\prime}\rvert$.
@@
Since the semi-circular curves ${SC}_{\kappa}^{j}$ and ${SC}_{\kappa}^{j+1}$, along with two other curves in the set $\{q(SC_{i,{\kappa}}):q\in\Gamma^{{N}}_{S_{\kappa}},$ $i=1,2,3,\text{ and }4\}$ bound a fundamental domain for $\Gamma_{S_{\kappa}}^{{N}}$, where $\Gamma_{S_{\kappa}}^{{N}}$ == ${h^{*}}_{(3,2)_{\kappa}},{h^{**}}_{(4,1)_{\kappa}}>$, we assert that ${h^{**}}_{(4,1)_{\kappa}}(SC_{\kappa}^{j})\cap SC_{\kappa}^{j+1}=\phi$. Let, $M$ and $N$ be two points lying on the semi-circle ${SC}_{\kappa}^{j}$ (where $j=1,2,3,...,\textbf{I}^{+}$). Note that the transformation ${h^{**}}_{(4,1)_{\kappa}}$ preserves any line through the points $\pm\tau$. Now, if we take a line ${H}_{(4,1)_{\kappa}}$ through the point $\{-\tau\}$ and the point $({x}_{j},0)$ we have the distance between the point $M$ and ${C}_{j}$ is $\{{r}_{j}-\textbf{(}\lvert{C}_{j}\rvert+\tau\textbf{)}\}$, which is away from the fixed point $\{-\tau\}$. Further, the distance between the point $N$ and ${C}_{j}$ is $\{{r}_{j}+\textbf{(}\lvert{C}_{j}\rvert+\tau\textbf{)}\}$. Also, the image of point $M$ under the transformation ${h^{**}}_{(4,1)_{\kappa}}$ situates on the line ${H}_{(4,1)_{\kappa}}$ and further from the point $\{-\tau\}$ than the point $N$.
@@
$\lvert{x}_{j+1}+\tau\rvert \displaystyle\leq$ $\{\lvert{x}_{j+1}\rvert+\lvert\tau\rvert\}$ (10)